A uniformly charge ring with radius R and having line charge density λ kept in xy plane

with its centre at origin. Find electric field intensity E, for this ring at a height L on z axis.


1
Expert's answer
2021-10-24T18:27:47-0400


We see that


"r^2=L^2+R^2,\\\space\\
\cos\theta=L/r=L/\sqrt{L^2+R^2}."

Determine the electric field intensity:


"\text dE_z=k·\text dq/r^2\cos\theta=\frac{kx\text dq}{(L^2+R^2)^{3/2}}.\\\space\\
E_z=\int \frac{kx\text dq}{(L^2+R^2)^{3/2}}=\frac{2\pi k\lambda RL}{(L^2+R^2)^{3/2}}."